OpenStudy (anonymous):

Solve the system 3x - 5y = 8 and 4x - 7y = 10. A. (1, -1) B. (2, 6) C. (6, 2) D. no solution E. infinitely many solutions

OpenStudy (lgbasallote):

3x - 5y = 8 4x - 7y = 10 -12x + 20y = -32 12x -21y = 30 -y = -2 y= 2 so x must be 6 via substitution so the o.p. is (6,2)